A project has annual cash flows of $6,500 for the next 10 years and then $11,000 each year for the following 10 years. The IRR of this 20-year project is 13.41%. If the firm's WACC is 12%, what is the project's NPV?
Answer #1
Step-1:Calculation of cost of project
IRR is the rate at which present value of cash inflows is equal to its cost.
Present value of annuity of 1 for 10 years = (1-(1+i)^-n)/i Where,
= (1-(1+0.1341)^-10)/0.1341 i = 13.41%
= 5.33847887 n = 10
Present value of 1 = (1+i)^-n
= (1+0.1341)^-10
= 0.28410998
Present value of cash flows of first 10 years = $   6,500.00 * 5.338479 = $ 34,700.11
Present value of cash flows of next 10 years = $ 11,000.00 * 5.338479 * 0.28411 = $ 16,683.87
Present value of cash flows of next 20 years $ 51,383.98
So, cost of project is $ 51,383.98
Step-2:Calculation of present value of project's cash inflows at WACC that is discounted rate
Present value of annuity of 1 for 10 years = (1-(1+i)^-n)/i Where,
= (1-(1+0.12)^-10)/0.12 i = 12.00%
= 5.65022303 n = 10
Present value of 1 = (1+i)^-n
= (1+0.12)^-10
= 0.32197324
Present value of cash flows of first 10 years = $   6,500.00 * 5.650223 = $ 36,726.45
Present value of cash flows of next 10 years = $ 11,000.00 * 5.650223 * 0.321973 = $ 20,011.43
Present value of cash flows of next 20 years $ 56,737.88
Step-3:Calculation of project's NPV
Present value of cash inflow $ 56,737.88
Less cost of project $ 51,383.98
Project's NPV $   5,353.90